Solve the exponential equation. Express the solution in exact form (ie. no calculator) then express the solution in
approximate form to two decimal places. 1) 2x + 8= 5
1
Expert's answer
2019-07-30T12:39:38-0400

"2^{x+8}=5".

"ln2^{x+8}=ln5".

"(x+8)ln2=ln5."

"x=\frac{ln5}{ln2}-8".

"x\approx-5.68."